OpenStudy (anonymous):

what is the midpoint if the endpoint are located at A(-5,7) and B(11,-9)

OpenStudy (jdoe0001):

\(\large { \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-5\quad ,&7)\quad &(11\quad ,&-9) \end{array}\qquad \left(\cfrac{x_2 + x_1}{2}\quad ,\quad \cfrac{y_2 + y_1}{2} \right)}\)

OpenStudy (anonymous):

m= (3,-1)